The weight of a can of stewed tomatoes is listed as 16 oz., but in actuality, the weight between cans may vary, and in fact is normally distributed with mean = 16.2 ounces and standard deviation = 0.1 oz. What is the probability that a consumer buys a can whose contents are actually less than the stated weight of 16 oz.?
Answer by stanbon(57967)

The weight of a can of stewed tomatoes is listed as 16 oz., but in actuality, the weight between cans may vary, and in fact is normally distributed with mean = 16.2 ounces and standard deviation = 0.1 oz. What is the probability that a consumer buys a can whose contents are actually less than the stated weight of 16 oz.?
----
z(16) = (16-16.2)/0.1 = -0.2/0.1 = -2
----
P(x < 16) = P(z < -2) = normalcdf(-1000,-2) = 0.0228
==============
Cheers,
Stan H.